# From Models to Network Topologies: A Topology Inference Attack in Decentralized Federated Learning

### v1

@@ Abstract
Federated Learning (FL) is widely recognized as a privacy-preserving machine learning paradigm due to its model-sharing mechanism that avoids direct data exchange. However, model training inevitably leaves exploitable traces that can be used to infer sensitive information. In Decentralized FL (DFL), the overlay topology significantly influences its models’ convergence, robustness, and security. This study explores the feasibility of inferring the overlay topology of DFL systems based solely on model behavior, introducing a novel Topology Inference Attack. A taxonomy of topology inference attacks is proposed, categorizing them by the attacker’s capabilities and knowledge. Practical attack strategies are developed for different scenarios, and quantitative experiments are conducted to identify key factors influencing the attack effectiveness. Experimental results demonstrate that analyzing only the public models of individual nodes can accurately infer the DFL topology, underscoring the risk of sensitive information leakage in DFL systems. This finding offers valuable insights for improving privacy preservation in decentralized learning environments.
@@ 1 Introduction
Federated Learning (FL) has emerged as a novel framework for enabling privacy-preserving Machine Learning (ML), facilitating collaborative model training among distributed clients without the necessity of sharing raw data McMahan et al. (2017). Conventional FL systems are predicated on a centralized architecture, referred to as Centralized FL (CFL), wherein a central entity is responsible for collecting, aggregating, and redistributing models to the clients. Nevertheless, this centralized architecture presents several challenges, such as processing bottlenecks and the single points of failure risk Beltrán et al. (2023). In response to these limitations, Decentralized FL (DFL) has been introduced, wherein model training and aggregation occur locally, and communication between nodes relies on P2P networks peer-to-peer (P2P) network. This approach eliminates dependency on a central server, mitigating risks associated with single points of failure Beltrán et al. (2024).
@@
Although FL, both centralized and decentralized, protect raw data privacy through its unique model-sharing mechanism, the training process inevitably leaves traces in the models. These traces may be susceptible to exploitation by malicious actors, potentially leading to inference attacks, including membership inference attacks, property inference attacks, and attribute inference attacks Hu et al. (2022). Previous studies have illustrated that such attacks pose a considerable risk to privacy within FL systems by elucidating sensitive information based on the behavior of the models Dayal et al. (2023). While much of the existing work focuses on the CFL, a notable research gap exists in exploring information leakage risks associated with DFL.
@@
Overlay network topology defines how participants are interconnected in an FL system. In CFL, the client-server architecture enforces a fixed, star-shaped topology. However, DFL leverages the P2P network, enabling flexible node connections to form diverse topologies. Existing studies demonstrate that topology significantly influences DFL models’ robustness and privacy-preserving capabilities Feng et al. (2024). Thus, topology should be considered sensitive data and a critical asset within DFL systems. However, theoretical analyses on how network topology affects model convergence remain sparse. Moreover, there is limited research investigating potential information leakage risks for this sensitive overlay topology data and strategies to safeguard this critical asset.
@@
This paper addresses this research gap by proposing a novel overlay topology inference attack targeting DFL systems. The proposed attack leverages the behavioral traces generated by models to uncover sensitive overlay topologies. The main contributions of this work are as follows: (i) the formulation of a taxonomy of topology inference attacks, classifying these attacks based on the attacker’s capabilities and knowledge; (ii) the development of practical strategies tailored to various attack types, accompanied by a quantitative analysis of critical factors that influence the efficiency of these attacks; (iii) rigorous experimental evaluations conducted across a range of datasets and topology configurations to empirically validate the proposed attack strategies; and (iv) the provision of insights aimed at informing the design of effective defensive mechanisms to safeguard sensitive information and preserve privacy within DFL systems.
@@ 2.1 Inference Attacks
Inference attacks exploit ML models to extract sensitive information without direct access to the underlying data Salem et al. (2018). By analyzing model behaviors or outputs, adversaries can infer properties of the data, model parameters, or system structure. As FL grows in adoption for its privacy-preserving capabilities, it also faces increased vulnerability to such attacks Yin et al. (2021). Inference attacks vary in objectives and strategies, with common types including:
  • Membership Inference Attack: Determines whether a specific data sample was part of the training dataset, often using shadow models or prediction confidence scores to differentiate between training and unseen data Shokri et al. (2017). This is critical in sensitive domains like healthcare or finance.
  • Model Inversion Attack: Reconstructs input data or infers its properties by iteratively optimizing inputs to match observed model outputs, posing risks in applications involving sensitive information like facial recognition or medical data Fredrikson et al. (2015).
  • Property Inference Attack: Infers characteristics of the training data, such as demographic distributions, by analyzing model updates or outputs. In FL, this can involve monitoring node-specific behaviors to deduce shared properties of local datasets Ganju et al. (2018).
  • Attribute Inference Attack: Targets specific attributes of training data, often by injecting malicious samples to influence the model’s learning process and make it vulnerable to data leakage Gong and Liu (2018).
These attacks not only compromise users’ sensitive information and undermine data security but also erode trust in FL’s privacy protection mechanisms. As a result, users may hesitate to participate in the training process, undermining the overall effectiveness of the FL system.
However, existing research on inference attacks in FL primarily focuses on CFL, with limited exploration of privacy leakage in DFL. Moreover, the critical role of DFL’s overlay topology has received insufficient attention in these studies.
@@ 2.2.1 Communication Systems
In communication systems, most protocols are designed to avoid ring formation, resulting in tree-like overlay topologies. Moreover, many topology inference studies commonly assume a single path between nodes Castro et al. (2004). A widely used approach involves leveraging ICMP-based protocols, such as the traceroute, to map network topologies effectively Jin et al. (2006). Tomography-based metrics, like packet loss rates, further aid in deducing a system’s topology: by observing loss rates from a source node to a target node, one can estimate the likelihood of a direct connection and thereby reconstruct the network Coates et al. (2002). Additionally, by measuring the timestamps of sent and received packets, it is possible to determine the number of forwarding hops, offering deeper insight into the overall network configuration Hou et al. (2020).
@@ 2.2.2 Social Networks
Social networks generally refer to the social structure formed by individuals in society connected by a specific relationship. In social networks, user interactions (e.g., likes, retweets) serve as valuable signals for topology modeling and inference by capturing how information diffuses across the network Gomez-Rodriguez et al. (2012). However, this behavioral data is often incomplete, especially in large-scale settings, necessitating methods like noisy sparse subspace clustering to infer network structures from partial observations Wang et al. (2016). Meanwhile, social networks are dynamic and time-sensitive, and they frequently exhibit small-world properties.Gong et al. (2012) incorporate these distributional patterns as prior knowledge in evolutionary models to better capture network changes over time.
@@ 3 Overlay Topology in DFL
The overlay topology in DFL significantly affects model convergence. The overlay topology is modeled as an undirected graph $G=(V,E)$, where $V$ is the set of nodes and $E$ represents the edges indicating direct communication links between nodes. The adjacency matrix $A$ encodes the graph structure, where:
@@
Thus, it can be seen that the convergence behavior of DFL is influenced by the topology captured by $P$. Initially, the aggregation dynamics (powers of $P$) dominate, reducing the influence of $M_{0}$. Over time, local updates drive the learning process, modulated by the decaying powers of $P$. This interplay underscores the importance of overlay topology in determining convergence efficiency and outcomes.
Given the critical role of overlay topology, it is essential to study potential information leakage through model updates and develop effective methods to protect sensitive topology information in DFL systems.
@@ 4 Problem Statement
[…]
@@
  • Scenario 1: The attacker knows all node models and all the datasets, and only partial edge information ($M^{\prime}=M,D^{\prime}=D,\emptyset\subsetneq E^{\prime}\subsetneq E$).
  • Scenario 2: The attacker knows all node models and none of the datasets, and only partial edge information ($M^{\prime}=M,D^{\prime}=\emptyset,\emptyset\subsetneq E^{\prime}\subsetneq E$).
  • Scenario 3: The attacker knows all node models and all the datasets, but no edge information ($M^{\prime}=M,D^{\prime}=D,E^{\prime}=\emptyset$).
  • Scenario 4: The attacker knows all node models and none of the datasets, but no edge information ($M^{\prime}=M,D^{\prime}=\emptyset,E^{\prime}=\emptyset$).
  • Scenario 5: The attacker knows partial models and none of the datasets, but has no knowledge of the edges ($\emptyset\subsetneq M^{\prime}\subsetneq M,D^{\prime}=\emptyset,E^{\prime}=\emptyset$).
@@ 5.1 Attack Metrics
In a DFL system, attackers may leverage the local data and local models available at each known node. Additionally, DFL’s decentralized communication mechanism allows attackers to access models transmitted from neighboring nodes. Thus, this paper proposes four attack metrics that attackers can exploit: relative loss, relative entropy, cosine similarity, and Euclidean similarity.
@@
    where $\mathcal{L}(f,D)$ is the loss function used by the model $f$ on dataset $D$.
@@
Relative loss and entropy require the attacker to use both the local model and the local dataset, which are suitable for attack scenarios 1 and 3. In contrast, cosine similarity and Euclidean similarity depend solely on the models and do not require access to the datasets, which are suitable for attack scenarios 2 and 4.
@@ 5.2 Attack Strategies
The attacker’s objective in a topology inference attack is to construct a predicted adjacency matrix $A^{\prime}$ of size $|V|\times|V|$ that closely approximates the ground-truth matrix. This can be viewed as a binary classification problem, determining whether each element $A^{\prime}_{ij}$ in $A^{\prime}$ equals one or zero.
Supervised Attack Strategy: In Attack Scenarios 1 and 2, since the attacker has knowledge of the true labels for some edges, a supervised classification approach is employed in this work to perform the attack. Given partial edges and their labels, these can be leveraged to train the attack model $M$. After training, $M$ predicts the unknown edges to generate the predicted adjacency matrix $A^{\prime}$, as shown in Algorithm 1.
Algorithm 1 Supervised Topology Inference Attack
1:$X$: Attack metrics, $Y^{k}$: Known edges, $Y^{m}$: Missing edges, $M$: Attack model
2:$A^{\prime}$: Predicted adjacency matrix
3:function TopologyInference($X,Y^{k},Y^{m}$)
4:     $D_{train}\leftarrow\textsc{ExtractNodePairsAndLabels}(X,Y^{k})$
5:     $D_{test}\leftarrow\textsc{ExtractNodePairs}(X,Y^{m})$
[…]
7:     $M.\text{train}(D_{train})$
8:     $Y^{\prime m}\leftarrow M.\text{predict}(D_{test})$
9:     $A^{\prime}\leftarrow\textsc{Combine}(Y^{k},Y^{\prime m})$
@@
11:function ExtractNodePairsAndLabels($X,Y^{k}$)
12:     $D_{train}\leftarrow[]$
13:     for each $(i,j)$ in $Y^{k}$ do
14:         $D_{train}.\text{append}(\text{GetNodeMetrics}(X[i],X[j]),\,Y^{k}[i,j])$
15:     return $D_{train}$
16:function ExtractNodePairs($X,Y^{m}$)
17:     $D_{test}\leftarrow[]$
18:     for each $(i,j)$ in $Y^{m}$ do
19:         $D_{test}.\text{append}(\text{GetNodeMetrics}(X[i],X[j]))$
20:     return $D_{test}$
Unsupervised Attack Strategy: For scenarios 3 and 4, where true labels are unavailable, this paper proposes an attack strategy based on unsupervised clustering, as shown in Algorithm 2. Since edges are known to fall into two categories, edge and non-edge group, the clustering algorithm $M$ is configured for two clusters ($K=2$). The attack metrics $X$ are then clustered by $M$, producing a predicted adjacency matrix $A^{\prime}$.
Algorithm 2 Unsupervised Topology Inference Attack
1:$X$: Attack metrics, $K=2$, $M$: Attack model
2:$A^{\prime}$: Predicted adjacency matrix
3:function TopologyInference($X$)
4:     $g_{1},g_{2}\leftarrow\textsc{ClusterNodeMetrics}(X,K)$
5:     $avg_{1}\leftarrow\textsc{CalculateMean}(g_{1})$
6:     $avg_{2}\leftarrow\textsc{CalculateMean}(g_{2})$
7:     if $avg_{1}<avg_{2}$ then
8:         $EdgeGroup\leftarrow g_{1},\;\;NonEdgeGroup\leftarrow g_{2}$
9:     else
10:         $EdgeGroup\leftarrow g_{2},\;\;NonEdgeGroup\leftarrow g_{1}$
11:     $A^{\prime}\leftarrow\textsc{CreateAdjacencyMatrix}(EdgeGroup)$
12:     return $A^{\prime}$
13:function ClusterNodeMetrics($X,K$)
14:     $M\leftarrow\text{InitializeClusteringModel}(K)$
15:     return $M.\text{fit}(X)$
16:function CreateAdjacencyMatrix($EdgeGroup$)
17:     Initialize $A^{\prime}$ as a zero matrix
18:     for each pair $(i,j)$ in $EdgeGroup$ do
19:         $A^{\prime}[i,j]\leftarrow 1$
20:     return $A^{\prime}$
These attack algorithms enable the implementation of diverse strategies, tailored to varying levels of attacker knowledge and capabilities, to effectively infer the topology of the DFL.

### v2

@@ Abstract
Federated Learning (FL) is widely recognized as a privacy-preserving machine learning paradigm due to its model-sharing mechanism that avoids direct data exchange. Nevertheless, model training leaves exploitable traces that can be used to infer sensitive information. In Decentralized FL (DFL), the topology, defining how participants are connected, plays a crucial role in shaping the model’s privacy, robustness, and convergence. However, the topology introduces an unexplored vulnerability: attackers can exploit it to infer participant relationships and launch targeted attacks. This work uncovers the hidden risks of DFL topologies by proposing a novel Topology Inference Attack that infers the topology solely from model behavior. A taxonomy of topology inference attacks is introduced, categorizing them by the attacker’s capabilities and knowledge. Practical attack strategies are designed for various scenarios, and experiments are conducted to identify key factors influencing attack success. The results demonstrate that analyzing only the model of each node can accurately infer the DFL topology, highlighting a critical privacy risk in DFL systems. These findings offer valuable insights for improving privacy preservation in DFL environments.
[…]
123
@@ 1 Introduction
Federated Learning (FL) has emerged as a novel framework for enabling privacy-preserving Machine Learning (ML), facilitating collaborative model training among distributed clients without the necessity of sharing raw data [25]. Conventional FL systems are predicated on a centralized architecture, referred to as Centralized FL (CFL), wherein a central entity is responsible for collecting, aggregating, and redistributing models to the clients. Nevertheless, this centralized architecture presents several challenges, such as processing bottlenecks and the single points of failure risk [3]. In response to these limitations, Decentralized FL (DFL) has been introduced, wherein model training and aggregation occur locally, and communication between nodes relies on a peer-to-peer (P2P) network. This approach eliminates dependency on a central server, mitigating risks associated with single points of failure [2].
@@
Although FL, both centralized and decentralized, protect raw data privacy through its unique model-sharing mechanism, the training process inevitably leaves traces in the models. These traces can be exploited by malicious actors, potentially leading to inference attacks, including membership inference attacks, property inference attacks, and attribute inference attacks [20]. Previous studies have illustrated that such attacks pose a considerable risk to privacy within FL systems by elucidating sensitive information based on the behavior of the models [7]. While much of the existing work focuses on CFL, a notable research gap exists in exploring information leakage in DFL.
@@
Overlay network topology defines how participants are interconnected in an FL system. In CFL, the client-server architecture enforces a fixed, star-shaped topology. However, DFL leverages the P2P network, enabling flexible node connections to form diverse topologies. Existing studies demonstrate that topology significantly influences DFL models’ robustness and privacy-preserving capabilities [10]. From a security perspective, exposure of the DFL network topology significantly facilitates malicious attacks. For example, in poisoning attack scenarios, topology awareness allows adversaries to pinpoint critical nodes, such as the central node in star topologies, thereby enabling them to amplify the poisoning intensity and enhance its propagation effectiveness [10]. Similarly, for network-level attacks, knowledge of key nodes within the DFL topology can lead to more targeted and effective Distributed Denial-of-Service (DDoS) attacks. Thus, topology information must be regarded as a critical asset, both in terms of optimizing DFL training performance and ensuring system security. However, theoretical analyses on how network topology affects model convergence remain sparse. Moreover, there is limited research investigating potential information leakage risks for this sensitive overlay topology data and strategies to safeguard this critical asset.
@@
Therefore, this paper improves this research gap by proposing a novel overlay topology inference attack targeting DFL systems. The proposed attack leverages the behavioral traces generated by models to uncover sensitive overlay topologies. The main contributions of this work are as follows: (i) the formulation of a novel taxonomy of topology inference attacks, classifying these attacks based on the attacker’s capabilities and knowledge; (ii) the development of practical strategies tailored to various attack types, accompanied by a quantitative analysis of critical factors that influence the efficiency of these attacks; (iii) rigorous experimental evaluations conducted across a range of datasets and real-world network topology configurations to empirically validate the proposed attack strategies; and (iv) the provision of insights aimed at informing the design of effective defensive mechanisms to safeguard sensitive information and preserve privacy within DFL systems.
@@ 2.1 Inference Attacks
Inference attacks exploit ML models to extract sensitive information without direct access to the underlying data [29]. By analyzing model behaviors or outputs, adversaries can infer properties of the data, model parameters, or system structure. As FL grows in adoption for its privacy-preserving capabilities, it also faces increased vulnerability to such attacks [36]. Inference attacks vary in objectives and strategies, with common types including:
  • Membership Inference Attack: Determines whether a specific data sample was part of the training dataset, often using shadow models or prediction confidence scores to differentiate between training and unseen data [31]. This is critical in sensitive domains like healthcare or finance.
  • Model Inversion Attack: Reconstructs input data or infers its properties by iteratively optimizing inputs to match observed model outputs, posing risks in applications involving sensitive information like facial recognition or medical data [11].
  • Property Inference Attack: Infers characteristics of the training data, such as demographic distributions, by analyzing model updates or outputs. In FL, this can involve monitoring node-specific behaviors to deduce shared properties of local datasets [12].
  • Attribute Inference Attack: Targets specific attributes of training data, often by injecting malicious samples to influence the model’s learning process and make it vulnerable to data leakage [14].
These attacks not only compromise users’ sensitive information and undermine data security but also erode trust in FL’s privacy protection mechanisms. As a result, users may hesitate to participate in the training process, undermining the overall effectiveness of the FL system. However, existing research on inference attacks in FL primarily focuses on CFL, with limited exploration of privacy leakage in DFL. Moreover, these studies have not adequately addressed the critical role of DFL’s overlay topology.
@@ 2.2.1 Communication Systems
In communication systems, most protocols are designed to avoid ring formation, resulting in tree-like overlay topologies. Moreover, many topology inference studies commonly assume a single path between nodes [5]. A widely used approach involves leveraging ICMP-based protocols, such as the traceroute, to map network topologies effectively [21]. Tomography-based metrics, like packet loss rates, further aid in deducing a system’s topology: by observing loss rates from a source node to a target node, one can estimate the likelihood of a direct connection and thereby reconstruct the network [6]. Additionally, by measuring the timestamps of sent and received packets, it is possible to determine the number of forwarding hops, offering deeper insight into the overall network configuration [19].
@@ 2.2.2 Social Networks
Social networks generally refer to the social structure formed by individuals in society connected by a specific relationship. In social networks, user interactions (e.g., likes, retweets) serve as valuable signals for topology modeling and inference by capturing how information diffuses across the network [13]. However, this behavioral data is often incomplete, especially in large-scale settings, necessitating methods like noisy sparse subspace clustering to infer network structures from partial observations [33]. Meanwhile, social networks are dynamic and time-sensitive, and they frequently exhibit small-world properties.[15] incorporate these distributional patterns as prior knowledge in evolutionary models to better capture network changes over time.
@@ 3 Overlay Topology in DFL
DFL is an ML paradigm that facilitates collaborative model training across multiple nodes without a central server. In DFL, each node updates its local model parameters using its private data and exchanges them with its directly connected neighbors. As the overlay topology in DFL significantly affects model convergence, this work presents the topology as an undirected graph $G=(V,E)$, where $V$ is the set of nodes and $E$ represents the edges indicating direct communication links between nodes. The adjacency matrix $A$ encodes the graph structure, where:
@@
Thus, it can be seen that the convergence behavior of DFL is influenced by the topology captured by $P$. Initially, the aggregation dynamics (powers of $P$) dominate, reducing the influence of $M_{0}$. Over time, local updates drive the learning process, modulated by the decaying powers of $P$. This interplay underscores the importance of overlay topology in determining convergence efficiency.
Given the critical role of overlay topology, it is essential to study potential information leakage through model updates and develop effective methods to protect sensitive topology information in DFL.
@@ 4 Problem Statement
Consider a DFL system represented by an undirected graph $G=(V,E)$, where $V$ is the set of participating nodes, and $E$ is the set of edges representing direct communication links between nodes. The adjacency matrix $A$ of size $|V|\times|V|$ encodes the network topology, as defined in Equation (1).
@@
  • Scenario 1: The attacker knows all node models and datasets, but only partial edge information ($M^{\prime}=M,D^{\prime}=D,\emptyset\subsetneq E^{\prime}\subsetneq E$).
  • Scenario 2: The attacker knows all node models, but none of the datasets, and only partial edge information ($M^{\prime}=M,D^{\prime}=\emptyset,\emptyset\subsetneq E^{\prime}\subsetneq E$).
  • Scenario 3: The attacker knows all node models and datasets, but does not have edge information ($M^{\prime}=M,D^{\prime}=D,E^{\prime}=\emptyset$).
  • Scenario 4: The attacker knows all node models, but neither datasets nor edge information ($M^{\prime}=M,D^{\prime}=\emptyset,E^{\prime}=\emptyset$).
  • Scenario 5: The attacker knows partial models, but neither datasets nor edge information ($\emptyset\subsetneq M^{\prime}\subsetneq M,D^{\prime}=\emptyset,E^{\prime}=\emptyset$).
@@ 5.1 Attack Metrics
In a DFL system, attackers may leverage the local data and local models available at each known node. Additionally, DFL’s decentralized communication mechanism allows attackers to access models transmitted from neighboring nodes. Thus, this paper proposes six attack metrics that attackers can exploit:
@@
    where $\mathcal{L}(f,D)$ is the loss function used by model $f$ on dataset $D$.
@@
  • Relative Sensitivity. This metric quantifies the sensitivity of model $f_{i}$’s predictions on dataset $D_{j}$ using the Jacobian norm. Relative sensitivity could capture the model’s susceptibility to small input perturbation.
    $\text{Relative Sensitivity}_{i,j}=\frac{1}{|D_{j}|}\sum_{x\in D_{j}}\left\|\frac{\partial f_{i}(x)}{\partial x}\right\|_{F}$ (10)
    where $\frac{\partial f_{i}(x)}{\partial x}$ is the Jacobian of $f_{i}$’s outputs with respect to input $x$, and $\|\cdot\|_{F}$ denotes the Frobenius norm.
@@
  • Curvature Divergence. This metric quantifies model divergence by analyzing implicit second-order information from consecutive parameter updates, reflecting curvature characteristics without explicit data dependence. Specifically, it compares parameter update vectors to capture similarity in model evolution indirectly:
    Curvature $\text{Curvature Divergence}(\Delta\mathbf{a},\Delta\mathbf{b})=\frac{\|\Delta\mathbf{a}-\Delta\mathbf{b}\|_{2}}{\frac{1}{2}(\|\Delta\mathbf{a}\|_{2}+\|\Delta\mathbf{b}\|_{2})}$ (13)
    where $\Delta\mathbf{a}=\mathbf{a}^{t}-\mathbf{a}^{t-1}$ and $\Delta\mathbf{b}=\mathbf{b}^{t}-\mathbf{b}^{t-1}$ denote parameter updates between rounds $t-1$ and $t$. A lower value indicates higher curvature divergence.
Relative metrics, including relative loss, entropy, and sensitivity, require access to both local models and local datasets, thus making them suitable for attack scenarios 1 and 3. While similarity metrics, i.e. cosine, Euclidean, and curvature similarity, depend only on the models and do not require dataset access, they are appropriate for attack scenarios 2 and 4.
@@ 5.2 Attack Strategies
In a topology inference attack, the attacker’s objective is to construct a predicted adjacency matrix $A^{\prime}$ that closely approximates the ground-truth adjacency matrix $A$.
5.2.1 Supervised Attack Strategy.
In attack scenarios 1 and 2, where the attacker has access to partial ground-truth edge labels, this work introduces EDGEPRE, a supervised edge classification framework for topology reconstruction. As shown in Algorithm 1, the model learns to infer the full adjacency matrix from node-level behavioral features and a small set of labeled node pairs.
Model Architecture.
Given $N$ nodes and their behavioral metrics $X\in\mathbb{R}^{N\times d}$, EDGEPRE directly uses an MLP-based edge classifier without any message passing. For each labeled pair $(i,j)$, the concatenated feature pair $[x_{i}\|x_{j}]$ is passed to a multilayer perceptron (MLP) decoder that outputs the predicted edge probability:
$\hat{y}_{ij}=\sigma(f_{\text{dec}}([x_{i}\|x_{j}])).$
Optionally, feature interaction terms such as $x_{i}*x_{j}$ and $|x_{i}-x_{j}|$ can be included to enhance the representation.
Training Objective.
The model is optimized using binary cross-entropy over the labeled edge set:
$\mathcal{L}=\sum_{(i,j,y_{ij})\in\mathcal{P}}\text{BCE}(\hat{y}_{ij},y_{ij}).$
Adjacency Reconstruction.
After training, EDGEPRE generalizes to all node pairs and produces a reconstructed soft adjacency matrix $A^{\prime}\in[0,1]^{N\times N}$, where each entry is computed via:
$A^{\prime}_{ij}=\sigma(f_{\text{dec}}([x_{i}\|x_{j}])).$
Algorithm 1 EDGEPRE (MLP-based)
1:Node features $X\in\mathbb{R}^{N\times d}$, labeled edge set $\mathcal{P}_{\text{train}}=\{(i,j,y_{ij})\}$, number of epochs $T$
2:Reconstructed adjacency matrix $A^{\prime}\in[0,1]^{N\times N}$
3:Initialize MLP decoder $f_{\text{dec}}$
4:for $t=1$ to $T$ do
5:     for each labeled pair $(i,j,y_{ij})\in\mathcal{P}_{\text{train}}$ do
6:         $h_{ij}\leftarrow[x_{i}\|x_{j}]$
7:         $\hat{y}_{ij}\leftarrow\sigma(f_{\text{dec}}(h_{ij}))$
8:     Compute loss: $\mathcal{L}\leftarrow\sum_{(i,j)}\text{BCE}(\hat{y}_{ij},y_{ij})$
9:     Update decoder parameters via gradient descent
10:Inference phase
11:for all $(i,j)$ where $1\leq i,j\leq N$ do
12:     $h_{ij}\leftarrow[x_{i}\|x_{j}]$
13:     $A^{\prime}_{ij}\leftarrow\sigma(f_{\text{dec}}(h_{ij}))$
14:return $A^{\prime}$
5.2.2 Unsupervised Attack Strategy.
For attack scenarios 3 and 4, where no edge labels are available, this work proposes INFERGAT, an unsupervised structure inference method based on a GAT encoder-decoder framework. The model predicts the latent adjacency purely from node-level behavioral metrics without requiring ground-truth edges.
Model Architecture.
The input $X\in\mathbb{R}^{N\times d}$ contains attack-related features such as similarity or entropy. A GAT encoder outputs embeddings $Z=f_{\text{enc}}(X)$, and a symmetric MLP decoder estimates pairwise edge strength via $A^{\prime}_{ij}=\sigma(\text{MLP}([z_{i}\|z_{j}]))$.
Training Objective.
The model is trained to reconstruct the behavioral structure with mean squared error:
$\mathcal{L}=\sum_{i,j}\left\|A^{\prime}_{ij}-X_{ij}\right\|^{2}$ (14)
The final output is the predicted adjacency matrix $A^{\prime}\in[0,1]^{N\times N}$.
Algorithm 2 INFERGAT
1:Attack metrics $X\in\mathbb{R}^{N\times d}$, learning rate $\eta$, epochs $T$
2:Predicted adjacency matrix $A^{\prime}\in[0,1]^{N\times N}$
3:Initialize GAT encoder $f_{\text{enc}}$ and decoder $f_{\text{dec}}$ with parameters $\theta$
4:for $t=1$ to $T$ do
5:     $Z\leftarrow f_{\text{enc}}(X)$triangleright Node embeddings via multi-head GAT
6:     $A^{\prime}_{ij}\leftarrow f_{\text{dec}}(z_{i},z_{j}),\quad\forall i,j$triangleright Decoder predicts edge weights
7:     $A^{\prime}\leftarrow\frac{1}{2}(A^{\prime}+A^{\prime\top})$triangleright Symmetrize adjacency
8:     $\mathcal{L}\leftarrow\sum_{i,j}\left\|A^{\prime}_{ij}-X_{ij}\right\|^{2}$triangleright Reconstruction loss
9:     $\theta\leftarrow\theta-\eta\cdot\nabla_{\theta}\mathcal{L}$triangleright Gradient descent update
@@
EDGEPRE and INFERGAT represent two attack strategies for topology inference in DFL. EDGEPRE leverages partial edge supervision to perform explicit edge classification, while INFERGAT relies solely on behavioral metrics to reconstruct structural patterns in a self-supervised manner. Together, they cover a broad range of adversarial capabilities and reflect realistic threat models across different attack scenarios.